Randy is fou
r times as old as Kenneth and Randy is 36 years older than Kenneth how old is Randy and kenneth

Answer by ankor@dixie-net.com(22740)

Randy is four times as old as Kenneth
r = 4k
and Randy is 36 years older than Kenneth
r = k + 36
how old is Randy and Kenneth
Replace r with 4k in the 2nd equation (from the 1st equation)
4k = k + 36
4k - k = 36
3k = 36
k = 36/3
k = 12 yrs is Kenneth's age
then obviously
r = 4*12
r = 48 yrs is Randy's age
:
Check in the 2nd equation
48 = 12 + 36